### v1

# How much do language models memorize?

Abstract
We propose a new method for estimating how much a model “knows” about a datapoint and use it to measure the capacity of modern language models. Prior studies of language model memorization have struggled to disentangle memorization from generalization. We formally separate memorization into two components: unintended memorization, the information a model contains about a specific dataset, and generalization, the information a model contains about the true data-generation process. When we completely eliminate generalization, we can compute the total memorization, which provides an estimate of model capacity: our measurements estimate that GPT-style models have a capacity of approximately 3.6 bits per parameter. We train language models on datasets of increasing size and observe that models memorize until their capacity fills, at which point “grokking” begins, and unintended memorization decreases as models begin to generalize. We train hundreds of transformer language models ranging from $500K$ to $1.5B$ parameters and produce a series of scaling laws relating model capacity and data size to membership inference.
Figure 1: Unintended memorization of uniform random data (Section 3). Memorization plateaus at the empirical capacity limit of different-sized models from the GPT-family, approximately 3.6 bits-per-parameter.
Figure 2: Unintended memorization of text across model and dataset sizes (Section 4). All quantities are calculated with respect to a large oracle model trained on the full data distribution.
1 Introduction
Figure 3: In our experiments on synthetic bitstrings, double descent occurs exactly when the dataset size begins to exceed the model’s capacity, when unintended memorization is no longer beneficial for lowering the loss.
Figure 4: Train and test losses of different model and dataset sizes trained on text. Double descent occurs when dataset size exceeds model capacity.
For the past several years, modern language models have been trained on increasingly large amounts of data, while parameter counts stay stagnant in the billions. For example, one recent state-of-the-art model (Dubey & et al, 2024) has 8 billion parameters (around $32GB$ on disk) but is trained on 15 trillion tokens (around $7TB$ on disk).
A long line of work (Carlini et al., 2019; Mireshghallah et al., 2022; Nasr et al., 2023; Zhang et al., 2023; Carlini et al., 2023b; Schwarzschild et al., 2024) questions whether such pretrained language models memorize their training data in a meaningful way. Most research approaches this problem either through the lens of extraction, aiming to recover full training data points from model weights, or membership inference, classifying whether a training point was present in the training data of a given model.
Studies of language model extraction argue that a data point is memorized if we can induce the model to generate it (Carlini et al., 2023b; Nasr et al., 2023; Schwarzschild et al., 2024). We argue that such generation does not necessarily serve as a proof of memorization. Language models can be coerced to output almost any string (Geiping et al., 2024); hence the fact that a model outputs something is not necessarily a sign of memorization. To address this issue, some researchers have suggested regularizing the input to the language model, such as by limiting its length Schwarzschild et al. (2024) or matching it to the prefix Carlini et al. (2023b) preceding the memorized sentence. However, even with these constraints, memorization cannot be conclusively proven, as the model’s ability to generalize may still be at play. For example, a good language model prompted to add two numbers can output the correct answer without having seen the equation before. In fact, a recent work Liu et al. (2025) shows that some of the instances than were previously thought as memorized do not even exist in the training set and their extractability is a result of generalization. Additionally, verbatim reproduction of a text is not a prerequisite for memorization; a model may still be memorizing specific patterns or sequences, such as every other token, without generating them verbatim.
Given that extraction is neither necessary nor sufficient, defining memorization accurately is a pressing concern. Existing mathematical definitions of memorization, such as those based on membership inference Shokri et al. (2017) and differential privacy Dwork (2006), are defined in the dataset/distribution level. This makes them inadequate for measuring memorization for certain instances (e.g. a particular textbook). Theoretical notions of ’influence’ Feldman (2020); Feldman & Zhang (2020) have been proposed to define memorization at the instance level, but they fall short of meeting our needs. These definitions focus on the capacity of a training algorithm to memorize inputs, whereas our interest lies in understanding how much a specific data point is memorized within a given model. This distinction is crucial, as we are concerned with the memorization properties of a single model, rather than the distribution of models generated by a particular training algorithm.
To bridge this gap, we propose a novel definition of memorization that quantifies the extent to which a model retains information about a specific datapoint. Our approach leverages the concept of compression rate in bits, where a model is considered to have memorized an input if it can be compressed to a significantly shorter encoding in the presence of the model. This framework draws inspiration from Kolmogorov information Kolmogorov (1963) theory and Shannon information Shannon (1948), but can be easily measured in practice using model likelihoods. We tackle the fundamental challenge of distinguishing between memorization and generalization (Prashanth et al., 2024) by decomposing memorization into two distinct components: unintended memorization, which captures the information a model stores about a particular dataset, and generalization, which represents the knowledge a model has acquired about the underlying data-generating process.
To understand our new quantities, we measure unintended memorization and generalization by training language models of varying capacity on datasets of different sizes. We first eliminate the question of generalization entirely by training on a dataset of random uniformly-sampled bitstrings. In this setting, we can exactly measure the amount of information contained about the data inside the model. This gives us a principled way to measure language model capacity when trained on uniform datasets of exact known information content. We find that GPT-style transformers can store between 3.5 and 4 bits of information in each model parameter, depending on model architecture and precision.
We then repeat our experiments with real text, where generalization is possible and even beneficial for learning. On real text, language models memorize up to a certain capacity, at which point they substitute unintended memorization for generalization, and begin to learn general, reusable patterns as opposed to sample-level specifics. Our framework shows that double descent phenomenon begins to occur at this point, when the data size exceeds the model capacity in bits.
Finally, we use our results to predict a scaling law for membership inference performance based on model capacity and dataset size. We show that membership inference follows a clean relationship based on model capacity and dataset size: bigger models can memorize more samples, and making datasets bigger makes membership inference harder. Our scaling laws extrapolate to larger models, and predict most modern language models are trained on too much data to do reliable membership inference on the average data point.
Figure 5: Bits memorized across training. This particular model is a GPT-style transformer with $6.86M$ parameters and a capacity of $23.9$ MB.
Figure 6: Capacity in bits-per-parameter for models trained on synthetic data. We estimate $\alpha=3.64$ bits-per-parameter for GPT models trained in half precision.
2 Memorization, intended and unintended
When a model $\theta=L(x)$ is trained using a training algorithm $L$ and a dataset $x\sim X$, some information is transferred from the sample $x$ to the model $\theta$. A key question in the memorization literature is determining how much of this stored information is intended versus unintended. In this work, we aim to provide a rigorous definition of memorization that satisfies certain properties:
  1. 1.
    Separation from generalization. Our notion of unintended memorization must be distinct from intended memorization, which we refer to as generalization. For example, consider a language model trained on the sample: Q: What is $2^{100}$? A: 1267650600228229401496703205376. When assessing how much of this training sample is memorized, we must account for the fact that performing simple math operations is expected from a language model.
  2. 2.
    Sample-level memorization. We need to define memorization for realizations of random variables, not the random variables themselves. Specifically, we want to determine how much unintended memorization of a sample $x$ occurs in a model $\theta$.
  3. 3.
    Independence from training algorithm. Our definition should be independent of the training algorithm $L$ and only a function of the final model $\theta$ and the sample $x$. This is crucial for language models, where we often only have access to the final model and target sample.
Previous works have attempted to define memorization for machine learning models. We aim to provide precise definitions of memorization that meet our criteria, and offer ways to measure it. See Appendix A.3 for a broader discussion on definitions of memorization.
2.1 A statistical view of memorization
Notation. In this section, we use capital letters (e.g. $X$, $\Theta$) to refer to random variables and lowercase letters to refer to instances of a random variable (e.g. $x\sim X$ and $\theta\sim\Theta$).
Information theory has developed well understood notions of information for random variables. For a random variable $X$, we often use $H(X)$, the entropy of $X$, to define the amount of information present in $X$. Moreover, for two distinct random variables $X,Y$, we can define $X\mid Y$ to be the uncertainty left in $X$ after fixing $Y$. Having defined this quantity, we can now measure mutual information between $X$ and $Y$ by subtracting the leftover information from the total information: $I(X,Y)=H(X)-H(X\mid Y)$.
Now assume we have a machine learning pipeline. We have a prior $\Theta$ on the underlying model that captures our dataset distribution $X$. And we have a learning algorithm $L$ that maps samples from $X$ to a trained model $\hat{\Theta}$. To understand how much information about $X$ is stored in $\hat{\Theta}$, we can use the notion of mutual information:
$\text{mem}(X,\hat{\Theta})$ $=I(X,\hat{\Theta})=H(X)-H(X\mid\hat{\Theta}).$
Note that this captures all the information about $X$ that is stored in $\hat{\Theta}$. As we discussed, we need our notion of memorization to account for generalization as well. So when measuring unintended memorization, we are only interested in the information that is present in $X\mid\Theta$, which is the uncertainty left in $X$ after fixing $\Theta$. Hence, we can define unintended memorization as
$\text{mem}_{U}(X,\hat{\Theta},\Theta)=I([X\mid\Theta],\hat{\Theta})=H(X\mid\Theta)-H(X\mid(\Theta,\hat{\Theta})).$
and then the generalization (or intended memorization) must be
$\text{mem}_{I}(\hat{\Theta},X,\Theta)=\text{mem}(X,\Theta)-\text{mem}_{U}(X,\hat{\Theta},\Theta)=I(X,\hat{\Theta})-I(X\mid\Theta,\hat{\Theta})$
Now that we have defined our notions of intended and unintended memorization we turn our attention to practically measuring them. Let us first state a proposition that enables measurement of unintended memorization:
Proposition 1 (Super-additivity of Unintended Memorization).
Assume $X=(X_{1},\dots,X_{n})$ is a dataset of $n$ i.i.d. samples. We have
$\sum_{i\in[n]}\text{mem}_{U}(X_{i},\hat{\Theta},\Theta)\leq\text{mem}_{U}(X,\hat{\Theta},\Theta)\leq H(\hat{\Theta}).$
This proposition shows that to measure a lower bound on the unintended memorization on the dataset level, we can sum per-sample memorization. On the other hand, the entropy of the information content of the trained model itself servers as an upper bound on the unintended memorization. Another implication of this implies that unintended memorization should scale with the dataset size but cannot exceed the total capacity of the model.
2.2 Measuring unintended memorization with Kolmogorov Complexity
Our definitions of memorization and generalization so far are defined using an “entropy-based" notion of information. This means our definitions can only be used for random variables. This brings big challenges in measuring memorization. All our variables in the definition of memorization are singletons. We have a single underlying model $\theta$, we have a single dataset $x=(x_{1},\dots,x_{n})$ and we have a single trained model $\hat{\theta}$ [^1] Note the switch to lowercase variables because we are now working with instances, not random variables.. It is impossible to measure the entropy (let alone conditional entropy) of the underlying variables using a single sample.
To this end, we switch to another notion of information based on compression, then later we show how this notion closely approximates the notion of memorization defined above. Kolmogorov complexity defines the information content of a string $x$, denoted as $H^{K}(x)$, to be the length of shortest representation of $x$ in a given computational model. Similarly, we can define the leftover information $x\mid\theta$, to be the shortest representation of $x$, when we have $\theta$ available as a reference. And the information content of $x\mid\theta$, denoted by $H^{K}(x\mid\theta)$, is the length of such description. Then, we can define mutual information in a similar fashion:
Definition 2 (Kolmogorov complexity).
Let $f$ be an arbitrary computational model that takes a set of inputs and returns an output (e.g. universal Turing machine). The shortest description of $x$ with respect to computational model $f$ is defined as $H^{K}(x)=\min_{f(p)=x}|p|.$ Also, the Kolmogorov complexity of $x$ relative to another string $\theta$ is defined as $H^{K}(x\mid\theta)=\min_{f(p,\theta)=x}|p|.$ And we define the Kolmogorov mutual information between $x$ and $\theta$ by $I^{K}(x,\theta)=H^{K}(x)-H^{K}(x\mid\theta).$ We assume inputs are bitstrings and $|p|$ is the bit length of the input.
Definition 3 (Kolmogorov memorization).
Let $\theta$ be a reference model that approximates the true distribution of data, and $\hat{\theta}$ be a model trained on a dataset $x=(x_{1},\dots,x_{n})$. For each $x_{i}$ we define the memorization of $x_{i}$ in $\hat{\theta}$ as $\text{mem}^{K}(\hat{\theta},x)=I^{K}(\hat{\theta},x).$ We also define intended and unintended variants of memorization:
$\text{mem}^{K}_{U}(x,\theta,\hat{\theta})=H^{K}(x\mid\theta)-H^{K}(x\mid(\theta,\hat{\theta}))\text{,\leavevmode\nobreak\ and\leavevmode\nobreak\ mem}^{K}_{I}(x,\theta,\hat{\theta})=\text{mem}^{K}(x,\hat{\theta})-\text{mem}^{K}_{U}(x,\theta,\hat{\theta}).$
There are known connections between Kolmogorov complexity and Shannon Entropy (Grunwald & Vitanyi, 2004). These results point at the conceptual connection between the two notions and imply that $\operatorname*{E}_{x\sim X}[H^{K}(x)]\approx H(X)$. Interestingly, this implies that our notion of Kolmogorov memorization closely approximates Shannon memorization.
Proposition 4.
Let $X=(X_{1},\dots,X_{n})$ be an i.i.d, dataset distribution parametrized by ground-truth model $\theta$. Let $L$ be a training algorithm mapping $X$ to $\hat{\Theta}$. Assume $H(\hat{\Theta})=\ell$ and $H(X_{i})=\ell^{\prime}$  [^2] The trained model and each data sample can be presented using $\ell$ and $\ell^{\prime}$ bits respectively. . Then we have $\Big{|}\operatorname*{E}_{\begin{subarray}{c}x\sim X\\ \hat{\theta}\sim L(x)\end{subarray}}\big{[}\text{mem}^{K}_{U}(x_{i},\hat{\theta},\theta)]\big{]}-\text{mem}_{U}(X_{i},\hat{\Theta},\theta)\Big{|}\leq\epsilon.$ for some constant $\epsilon$ independent of $\theta,\ell,\ell^{\prime}$ and $n$. Moreover, we have tail bounds
$\Pr_{\begin{subarray}{c}x\sim X\\ \hat{\theta}\sim L(x)\end{subarray}}\Big{[}\big{|}\Gamma(x,\hat{\theta})\big{|}\geq r+c\Big{]}\leq e^{-2\frac{c^{2}}{\ell\ell^{\prime}}}.$
2.3 Estimating Kolmogorov with likelihoods
Fixing our notion of Kolmogorov memorization, we now describe how we can estimate $H^{K}$ in different setups. Note that exact calculation of Kolmogorov complexity is known to be uncomputable (the decision version of is undecidable). However, we can still approximate it using the best available compression schemes. Below, we summarize how we approximate each term in our definition.
  • $H^{K}(x\mid\hat{\theta})$: Here, $\hat{\theta}$ is the trained target model, which does not necessarily capture the true data distribution. Because compression rate is inherently tied to the likelihood under a predictive model model (Shannon, 1950), we can easily estimate $H^{K}(x\mid\hat{\theta})$ using $p(x\mid\hat{theta})$, the likelihood of $x$ under the target model.
  • $H^{K}(x\mid\hat{\theta},\theta)$: In this case, the compression algorithm has access to both target and reference models. We simply compute $\max\{p(x\mid\hat{\theta},p(x\mid\theta)\}$. In practice, our choice of reference model is a larger model with the same architecture as $\theta$ trained for many steps on a much wider data distribution.
$\text{Capacity}(\theta)$    [bits] $\alpha$    [bpp]
$n_{\text{layer}}$ $d_{\text{model}}$ Params fp32 bf16 fp32 bf16
1 32 8.04×$\times 10^{4}$ 3.39×$\times 10^{5}$ 3.16×$\times 10^{5}$ 4.23 3.93
64 1.85×$\times 10^{5}$ 7.27×$\times 10^{5}$ 6.93×$\times 10^{5}$ 3.92 3.74
128 4.69×$\times 10^{5}$ 1.71×$\times 10^{6}$ 1.69×$\times 10^{6}$ 3.65 3.61
256 1.33×$\times 10^{6}$ 4.15×$\times 10^{6}$ 3.83×$\times 10^{6}$ 3.12 2.88
2 32 9.31×$\times 10^{4}$ 3.87×$\times 10^{5}$ 3.31×$\times 10^{5}$ 4.16 3.56
64 2.35×$\times 10^{5}$ 9.60×$\times 10^{5}$ 9.27×$\times 10^{5}$ 4.08 3.94
128 6.67×$\times 10^{5}$ 2.66×$\times 10^{6}$ 2.60×$\times 10^{6}$ 3.99 3.89
256 2.12×$\times 10^{6}$ 8.49×$\times 10^{6}$ 7.76×$\times 10^{6}$ 4.01 3.66
4 32 1.18×$\times 10^{5}$ 4.65×$\times 10^{5}$ 3.99×$\times 10^{5}$ 3.92 3.37
64 3.35×$\times 10^{5}$ 1.34×$\times 10^{6}$ 1.14×$\times 10^{6}$ 3.98 3.39
128 $\times 10^{6}$ 4.02×$\times 10^{6}$ 3.75×$\times 10^{6}$ 3.78 3.53
256 3.70×$\times 10^{6}$ 1.36×$\times 10^{7}$ 1.30×$\times 10^{7}$ 3.68 3.51
8 32 1.69×$\times 10^{5}$ 5.12×$\times 10^{5}$ 4.85×$\times 10^{5}$ 3.02 2.86
64 5.35×$\times 10^{5}$ 2.05×$\times 10^{6}$ 1.71×$\times 10^{6}$ 3.83 3.19
128 1.86×$\times 10^{6}$ 7.23×$\times 10^{6}$ 6.49×$\times 10^{6}$ 3.89 3.49
256 6.86×$\times 10^{6}$ 2.71×$\times 10^{7}$ 2.51×$\times 10^{7}$ 3.96 3.65
Mean (±$\pm 0.1$):
3.83 3.51
Figure 7: Model capacity estimates across different widths and depths in full and half-precision. Doubling precision from bfloat16 to float32 only increases model capacity from $3.51$ to $3.83$ bits-per-parameter.
Figure 8: Scaling law curves for membership inference overlaid with empirical data, shown in circles.
3 Model Capacity for Memorization
Unintended memorization provides us a principled way of measuring the precise number of bits a model $\theta$ knows about a datapoint $x$. If we add up the information for each datapoint in a dataset, we can measure the total amount of bits a model knows about the dataset. And in cases where generalization is not possible because each datapoint is completely independent, we can estimate the capacity of a given model $\theta$ by summing per-datapoint unintended memorization.
3.1 Defining model capacity
We first formalize this notion of memorization capacity for a particular language model $\theta$. Capacity is the total amount of memorization that can be stored in $\theta$ across all its parameters.
Definition 5 (Capacity).
Let $X$ be a distribution and $L\colon X\to\Theta$ a learning algorithm. We define the capacity of the learning algorithm $L$ to be
$\text{Capacity}(L)$ $=\max_{X}\text{mem}\big{(}X,L(X)\big{)}$
When the model capacity is reached, $\text{mem}(X,L(X))$ will no longer increase with dataset size. In practice, we can compute capacity by training to saturation on varying sizes of $X$ and computing the maximum memorization.
3.2 Measuring model capacity with synthetic sequences
In this section we measure the capacity of Transformer language models. Our goal is to instantiate multiple datasets and distributions and measure the memorization of them when training a single model $\theta$. Then, we take the maximum over all datasets to approximate of the model’s capacity. For instantiating our datasets, each token is uniformly sampled from a predefined set of tokens independent of the previous tokens.
To approximate $H^{k}(x\mid\theta,\hat{\theta})$, we can directly compute entropy under the trained model to calculate the shortest description of the dataset conditioning on $\hat{\theta}$. Subtracting the two, we can approximate the unintended memorization $\text{mem}_{U}(X,L(X))$. Since the process for sampling the data is completely random, there is no generalization to be stored within $\hat{\theta}$ (that is, $\text{mem}^{U}(X,L(X))\approx\text{mem}(X,L(X))$).
Observe that when we sample synthetic sequences from a uniform distribution, we can compute their Shannon information exactly. Given a dataset size $N$, we construct a dataset of $N$ sequences, each of $S$ tokens. Given a vocabulary size $V$, we can calculate the total entropy of a dataset $x^{i}$ with such parameters by $H(x^{i})=NS\log_{2}V$. Then we calculate the compressed form $x^{i}$ using entropy under $\hat{\theta_{i}}$ to compute the code length and use this as an approximation of $H^{K}(x^{i}\mid\hat{\theta_{j}})$. Then we calculate the $\text{mem}(x^{i},\hat{\theta_{i}})=H(x^{i})-H^{K}(x^{i}\mid\hat{\theta_{j}})$ and compute a model’s capacity as the maximum amount of memorization over all datasets.
Experimental details.
In accordance with Kaplan et al. (2020), we train models with the GPT-2 architecture (Radford et al., 2019) initialized from scratch. Our models have between $1$ and $8$ layers, hidden dimensions scaled from $32$ to $512$, and from $100K$ to $20M$ parameters. We train models for $10^{6}$ steps with a batch size of $2048$. We use the Adam optimizer. All models are trained on a single A100 GPU in bfloat16 precision, and we use gradient accumulation if a batch cannot fit in memory. Unless otherwise noted, we set vocabulary size $V=2048$, sequence length $S=64$ and vary only the number of points in a dataset. We train each model on each dataset size over five random seeds, which affect both model initialization and the dataset sampling.
Results.
We plot memorization across model and data sizes in Figure 2. This allows us to visualize unintended memorization amounts (y-axis) across dataset sizes (x-axis) grouped by model size (line color). We observe a striking plateau once a model reaches its capacity. Given the dataset is large enough, models exhibit an upper bound in net memorization, regardless of data size. Small datasets are completely memorized by all models with enough capacity.
We estimate the capacity of each model as the maximum amount of unintended memorization in bits measured across all dataset sizes. We then compare this capacity to the model size in Figure 6. Interestingly, even at this small scale, we see a very smooth relationship between observed capacity (maximum memorization measured over all datasets) and model parameters. We plot this relationship in Figure 6: under these settings, our models consistently memorize between $3.5$ and $3.6$ bits per parameter. This corroborates the findings of prior work such as (Roberts et al., 2020; Lu et al., 2024), which noticed that fact storage scales linearly with model capacity. Ours is a slightly larger estimate than Allen-Zhu & Li (2024), which estimated via quantization that models can store around $2$ bits per parameter.
Since our models are learned via gradient descent, they are not guaranteed to find the global optima; thus, we are only ever measuring a lower bound on model capacity. We take a closer look at the training curves to analyze the convergence of our 8M parameter language model. We plot model convergence throughout training in Figure 6.
In this case, all datasets from 16,000 to 4M samples fall within a range of $3.56-3.65\times 10^{6}$ bits memorized. This indicates that our measurements are robust within an order of magnitude, and we do not expect to memorize significantly more information by training for more steps. This finding also confirms our hypothesis that capacity scales roughly with parameter count. The two largest datasets (4M and 8M samples, respectively) converge to total memorization of $2.95\times 10^{6}$ and $1.98\times 10^{6}$ bits memorized. We expect that their memorization rates would continue to increase toward the capacity had we trained for more epochs.
How does precision affect capacity?
One natural question is how our estimates for $\alpha$ depend on the precision of language model training. In fact, although most software defaults to training in 32-bit precision, recent work has shown that language models can be quantized to fewer than 2 bits per parameter and still retain much of their utility. Since all other experiments have been conducted in bfloat16 precision, rerun our experiments in full fp32 precision to analyze the effect on capacity. Across model sizes, we observe a small increase in capacity, and an increase in $\alpha$ from 3.51 to $3.83$ bits-per-parameter on average. This is far less than the actual 2x increase in the bits of $\theta$, indicating that most of the extra model bits added when increasing precision from bfloat16 to float32 are not used for raw storage.
4 Disentangling Unintended Memorization from Generalization
Our previous experiments analyzed the memorization and membership inference properties of synthetic bitstrings. We now turn to measuring memorization of text. Unlike randomly generated sequences, learning from text data is a mix of both unintended memorization (sample-level) and generalization (population-level). Therefore, as a reference model, we use the model of an equal parameter count trained on the maximum amount of data (in this case, the entire dataset). [^3] Restricting the computational power of the reference model relates its predictions to $\mathcal{V}$-information (Xu et al., 2020) which measures the “usable” information available in a signal, when accounting for model size. We also consider an oracle reference model, which is the model that achieves the best compression rate (lowest loss) on the evaluation dataset, and may have many more parameters.
Experimental details.
We repeat the experiments from Section 3.2, substituting our synthetic datapoints for real text. To obtain a distribution of real-world text data, we could use any pre-training scale text dataset; we use the recently proposed FineWeb dataset (Penedo et al., 2024) as it follows state-of-the-art deduplication practices. We use sequences of $64$ tokens but perform an additional deduplication step to ensure perfect deduplication (otherwise, that $1-2\%$ of sequences become duplicates when truncating to $64$ tokens). We find careful deduplication extremely important for faithfully measuring extraction rates. As in the previous subsection, we pretrain models of varying sizes on different-sized text datasets and measure the unintended memorization of each model-dataset pair. In addition to memorization, we measure membership inference performance according to a standard loss-based membership inference procedure; we also compute exact extraction rates by greedily decoding prefixes of different lengths.
Results.
We first observe that the sample-level unintended memorization increases with model parameters and decreases with training set size (Figure 4). When we measure unintended memorization with respect to an oracle reference model (Figure 2), memorization steadily increases as our smaller model is able to learn more about the small training set than the oracle, and then decreases as our model starts to generalize and perform on average worse than the (higher-capacity) oracle.
We then compute unintended memorization across the entire dataset with respect to the restricted model (Figure LABEL:fig:text-km-total).
Dataset-to-capacity ratio predicts double descent.
We observe from the train and test loss that for larger datasets the model only begins to generalize (i.e. evaluation loss decreases) once its capacity is reached, which takes approximately $10^{5}$ samples, depending on parameter count. As in Nakkiran et al. (2019) we plot the ratio between the dataset size and model capacity (Figure 4). Unlike prior work, in our experiments we can compute the exact dataset size (based on the compression rates of the reference model) and exact model capacity (based on our estimate of $\alpha$).
We clearly observe double descent evaluation performance decreases as the training set size nears model capacity, and then rapidly drops as the dataset capacity exceeds the capacity of the model. Our observations offer an intuitive explanation for double descent (Belkin et al., 2019; Nakkiran et al., 2019): double descent begins exactly when the data capacity exceeds the model capacity. One theory is that once the model can no longer memorize datapoints individually, it is forced to share information between datapoints to save capacity, which leads to generalization.
Figure 9: Extraction rates of 64-token training sequences across prefix lengths, for both train and evaluation.
Figure 10: Membership inference F1 across dataset sizes. In this case, F1 score of 0.5 implies random guessing.
Figure 11: Membership inference vs 32-token-prefix suffix extraction rate. Membership inference is generally easier than extraction.
Generalization explains nonzero extraction rates.
We measure extraction rates on the full training set and 10,000 non-overlapping test samples (Figure 18). We note that for 32-token prefixes, 100% are extractable for very small training set sizes; predictably, all extraction numbers decrease with training set size. When the dataset sizes grows sufficiently large, the extraction rate does not go fully to zero; however, it converges to nearly exactly the test extraction rate. In other words, when our (deduplicated) dataset grows sufficiently large, all successful training data extraction is attributable to generalization.
5 Memorization and Membership
Our training settings allow total control over the train and test data and come with perfect deduplication. This makes our setting ideal for studying the relationship between model size, dataset size, and membership inference success rate.
All of our membership inference results come from a standard loss-based membership inference (Yeom et al., 2018; Sablayrolles et al., 2019). The method is very simple: we set a cutoff loss value to predict whether a sample is or is not a member of the training dataset.
5.1 Membership in synthetic and text data
Synthetic data.
For each of our models trained on synthetic data, we plot the success rate of the membership inference attack attack across dataset sizes. We show results in Figure 15. Above a certain dataset size, membership inference starts to fail in the average case. This finding indicates that if the dataset size is too large compared to the model, membership inference of an average training sample may not be possible.
Text.
For each of our models trained on text, we use unused non-overlapping data from FineWeb to perform a standard loss-based membership inference (Yeom et al., 2018; Sablayrolles et al., 2019) on each model and plot performance across dataset sizes (11). For a fixed model size, membership inference gets more difficult as the size of the data increases. When comparing membership inference to extraction (Figure 11), membership inference is strictly higher in every case; in some cases we can infer training dataset membership quite well (score of $0.97$) with an extraction rate of 00.
5.2 Scaling laws for Membership
In this section we develop a set of predictive models for memorization. Specifically, we predict the F1 score of a loss-based membership attack given token count, number of examples, and model parameter count. We then validate our predictions on models from $500$K to $1.5$B parameters.
5.2.1 Functional forms
We observe that for a fixed model capacity, membership inference follows a roughly sigmoidal form with respect to dataset size. The intuitive explanation is that M.I. is easy for large models overfit to tiny datasets, so its score begin at 1; as dataset size increases, differentiating train from test data by loss becomes more and more difficult, eventually decaying toward 0.5.
We reuse the data collected in our text experiments (Section 4) to solve for constants $c_{1},c_{2},c_{3}$ in the following equation:
$\text{Membership}_{F_{1}}(\theta,\mathcal{D})=\dfrac{1}{2}{(1+c_{1}\sigma(c_{2}(\dfrac{\text{Capacity}(\theta)}{|\mathcal{D}|}+c_{3}))}$
where $\sigma(x)=\frac{1}{1+e^{-x}}$.
Limiting behavior.
We observe that as $|\mathcal{D}|\rightarrow\infty$, performance of our membership inference attack decreases to $0.5$ (essentially random performance). For a model trained on an infinite dataset, our law predicts both membership inference and extraction to be impossible.
Fitting.
We use a non-linear least squares solver to find optimal values for $c_{1},c_{2},c_{3}$. Solutions found are $c_{1}=1.34$, $c_{2}=-0.034$, and $-33.14$. We plot the scaling laws along with observed data in Figure LABEL:fig:memb-inf-heatmap. Although the sigmoidal function is slightly simplistic (the points do not perfectly fit) our fit produces estimates within $1-2\%$ of observations.
5.2.2 Validation on larger models
We note that all contemporary language models trained with a tokens-per-parameter ratio of $10^{2}$ or higher, which according to our laws would imply membership inference score of $0.5$ – that is, within our formulation, statistically significant loss-based membership inference is not possible.
To validate our predictions, we train models with expected membership $F1$ scores of 0.55, 0.75, and 0.95. For model sizes we select GPT-2 small ($125$M params) and GPT-2 XL ($1.5$B params). Using our scaling law, we solve for the dataset size required to get the desired membership inference score for the given model size (see Table 1 for more information). We train models on the estimated dataset size and measure F1 score (Figure LABEL:fig:memb-inf-heatmap). Our predictions are generally within $1.5$ points of the true F1 score; the score is most inaccurate for estimated F1 of 0.75, which is the point where the sigmoid is steepest. In general, the accuracy of our results indicates that our empirical model of membership inference is relatively accurate and provides evidence for why membership inference attacks fail on models trained on extremely large datasets (Das et al., 2024; Duan et al., 2024; Maini et al., 2024).
6 Conclusion
We propose a new definition of memorization that allows us to measure the exact number of bits a model knows about a dataset. We use our definition to measure the capacity of modern transformer language models and analyze how measurements such as extraction and F1 score scale with model and dataset size. We also propose a scaling law for membership inference and validate it on larger models. Our results help further practitioner understanding of how language models memorize and what they might (or might not) be memorizing across model and dataset scales.
7 Acknowledgements
Thanks to the many folks who helped us improve our paper, including Karen Ullrich, Niloofar Mireshghallah, Mark Ibrahim, Preetum Nakkiran, and Léon Bottou.
Appendix A Appendix
A.1 Related Work
Language models and compression.
Shannon’s source coding theorem (Shannon, 1948) first formalized the duality between prediction and compression. The connection between language modeling and compression was studied as far back as Shannon (1950), which observed that more accurate models of English can compress text in fewer bits. Other works note the connection between Kolmogorov complexity (Kolmogorov, 1965) and Shannon information in detail (Grunwald & Vitanyi, 2004). Delétang et al. (2024) investigate using modern transformer-based language models as compressors. We use compression as a tool to measure memorization in models.
Language model capacity.
(Arpit et al., 2017) formalize the idea of effective capacity of a model and its training procedure; they also observe that both representation capacity and training time have a strong impact on empirical model capacity. Several other works measure language model capacity in the number of facts or random labels that can be memorized by a network such as an RNN (Collins et al., 2017; Boo et al., 2019) or transformer (Roberts et al., 2020; Heinzerling & Inui, 2021; Allen-Zhu & Li, 2024), sometimes under quantization. A few research efforts (Yun et al., 2019; Curth et al., 2023; Mahdavi et al., 2024; Kajitsuka & Sato, 2024) have developed theoretical estimates for the capacity of different model architectures, although none have yet scaled to multi-layer modern transformers. We are the first to measure a clear upper-bound in model capacity.
Alternative definitions of memorization.
Unintended memorization is deeply related to the many other definitions of memorization proposed in the literature. We provide a detailed comparison in Table LABEL:tab:defns and Section A.3.
A.2 Related Work: Definitions of Memorization
Prior definitions of memorization.
Carlini et al. (2019) defined a string $m$ as memorized by a language model $\theta$ if the second half of $m$ can be generated greedily when prompting the model with the first half. Following this, Nasr et al. (2023) introduced extractable memorization, where model $\theta$ is said to memorize $m$ if an adversarial prompt $p$ can be found that generates $m$. Mireshghallah et al. (2022) and Schwarzschild et al. (2024) refined this definition by restricting $p$ to a certain number of tokens, preventing it from containing the entire $m$. However, even this definition has limitations: for example, generating the sequence “cat cat cat … cat" with the prompt "repeat cat 1000 times" does not necessarily indicate memorization. Carlini et al. (2019) use perplexity or likelihood, one measure of the compressibility of a sequence, in an effort to distinguish highly memorized sequences from merely easy-to-compress ones. One additional definition of note is counterfactual memorization (Zhang et al., 2023), which measures the impact of a single datapoint on training; this can be seen as an instantiation of our definition where a different model of the same family is used as a reference model. Overall, all these works regarded memorization in terms that can be seen as forms of compression, although did not explicitly define it as such.
Finally, a concurrent work (Cohen et al., 2024) proposes a theoretical definition for memorization also relying on Kolmogorov.
Some of our findings also relate to the discovery of double descent in machine learning (Belkin et al., 2019; Nakkiran et al., 2019) and language modeling (Xia et al., 2023), as well as general discussions of memorization and generalization in deep learning (Zhang et al., 2017; Tänzer et al., 2022).
Here, we discuss other definitions of memorization.
A.3 Other notions of memorization
In this section we list multiple other notions of memorization and compare it with our definition. We specifically focus on why these notions do not satisfy all of our requirements.
  • Stability-based notions of memorization. There are notions of privacy and memorization that deal with “stability” of the training algorithm to small changes in the training set. Most notably, differential privacy Dwork (2006) considers the worst-cast drift of the model distribution when a single data point changes. Another notion of memorization in Feldman (2020) is based on the change of the model prediction on a point $x$, when we add the labeled pair $(x,y)$ to the training set of a classification/regression model. Both of these notions are crucially relying on the learning algorithm and how it behaves. Moreover, the definition of differential privacy is not ideal for our case because it is a worst-case definition and cannot be applied at sample/model level. While the notion of memorization in Feldman (2020) does not have this particular issue, it suffers from the fact that it only applies to classification models and mostly deals with the memorization of the association between the label ($y$) and input ($x$), and not the memorization of $x$ itself. These issues make these notions not ideal for our case.
  • Extraction-based memorization. There are multiple works in the literature (Carlini et al., 2019; Mireshghallah et al., 2022; Nasr et al., 2023; Zhang et al., 2023; Carlini et al., 2023b; Schwarzschild et al., 2024) that define memorization of samples in language models based on how easy it is to extract that sample. Specifically, when trying to understand the extent of memorization of a sample $x$ in a model $\theta$ they measure some notion of complexity for the task of eliciting the model to output $x$. Although these notions are great in that they only take a model $\theta$ and a sample $x$, they still do not account for generalization. Considering our running example of the following training sample: "What is $2^{100}$? (A: $1,267,650,600,228,229,401,496,703,205,376$)", this will be identified as highly memorized by almost all of the extraction based notions of memorization. Another issue with these definitions are that they are heavily dependent on the details of decoding algorithm. This is not ideal as we do not expect the memorization of a sample $x$ in a model $\theta$ to depend on the detailed parameters we use to generate samples using $\theta$.
    The work of Schwarzschild et al. (2024) in this category is the closest to ours. This work which is based on prompt-optimization, optimizes a short prompt $p$ to make the model elicit $x$, then it calls the sample $x$ memorized, if length of $p$ is less than $x$. Although this definition is close to our definition in using compression, it still does not account for generalization of the model. Moreover, it focuses on a specific way of compression through prompting. We posit that compression through prompting is an inferior compression scheme and can often lead to compression rates greater than 1.
  • Membership/attribute inference. Membership inference Shokri et al. (2017) and attribute inference attacks Jayaraman & Evans (2022) have been used for empirically measuring the privacy of machine learning algorithms. These notions which usually aim at approximating the stability notions of memorization are suffering from the same shortcomings. They rely heavily on the learning algorithm and the data distribution. Moreover, they fail at providing a sample level notion of memorization. For example, the obtained accuracy for membership inference attack is only meaningful in the population level. This is because various attack may have different true positives for membership, and the union of all these true positive across different attack may cover the entire training set, rendering it unusable as a sample level notion of memorization.
  • Data copying in generative models. There are some interesting notions of memorization designed specifically for generative modeling where a generative model may output a certain portion of training samples (Bhattacharjee et al., 2023; Carlini et al., 2023a). These notions are similar to extraction based definition of memorization but they are more lenient in that they only require extraction of part of the training data. However, they still suffer from the same challenges as of extraction based definitions.
A.4 Compression with language models beyond arithmetic coding
Shannon (1948) noted that the optimal compression method for a given source is one that assigns codes to symbols such that the average code length approaches the entropy of the source. Arithmetic coding (Pasco, 1977; Rissanen, 1976) is known to be one optimal way to compress text given a distribution over symbols; it was used in (Delétang et al., 2024) to compress text using modern language models.
Although arithmetic coding is known to be optimal for samples generated from the random process of choice, it may still be sub-optimal for cases where the compressed samples are correlated with the choice of random process. Specifically, in language modeling, the training data is highly correlated with the model itself and hence we might need to treat them differently. For instance, we know from previous work that the models behavior on training data points is different from random samples. A large portion of training data can be generated using greedy decoding (Carlini et al., 2023b; Liu et al., 2025) which is a behavior not expected for randomly sampled data. To this end, we design a new compression technique, a generalization of arithmetic coding.
Ensemble compression.
Sampling from language models involve two key parameters $k$ for $top_{k}$ selection and $t$ for temperature. We design a compression method that sets these parameters adaptively. For instance, for cases where we know we can decode the next $100$ tokens in a greedy fashion, we set $k=1$ to reduce the bit length of arithmetic code. Changing the setup of the coding scheme itself requires a new token to be injected and wastes some number of bits, but it could still be beneficial for the code length. Our compression program uses dynamic programming to find the optimal code with injection of these new tokens in the middle of the text. Notably, our algorithm runs in time $O(n*T)$, where $n$ is the number of tokens and $T$ is the number of possible setups (combination of $t$ and $k$) that we allow.
A.5 How reliable are our linear estimates of capacity?
$d_{emb}$ $n_{layer}$ $|\theta|$ $|D|$ Predicted F1 Observed F1
GPT2-XL 1600 48 1,556,075,200 170,654,583 0.55 $54.61\pm 1.3$
76,795,021 0.75 $71.08\pm 0.4$
18,851,574 0.95 $95.85\pm 0.8$
GPT2-Medium 768 12 123,702,528 13,566,442 0.55 $53.44\pm 1.1$
6,104,935 0.75 $65.69\pm 0.6$
1,498,634 0.95 $97.98\pm 0.3$
Table 1: Dataset sizes that our scaling law predicts will produce a given membership inference F1, along with empirical values.
$S$ Params. Memorized Expected Error
4 $6.59\times 10^{5}$ $1.73\times 10^{5}$ $1.80\times 10^{5}$ 4.19
8 $6.60\times 10^{5}$ $3.54\times 10^{5}$ $3.60\times 10^{5}$ 1.80
16 $6.61\times 10^{5}$ $7.15\times 10^{5}$ $7.21\times 10^{5}$ 0.84
32 $6.63\times 10^{5}$ $1.44\times 10^{6}$ $1.44\times 10^{6}$ 0.41
64 $6.67\times 10^{5}$ $2.29\times 10^{6}$ $2.36\times 10^{6}$ 2.97
128 $6.75\times 10^{5}$ $2.36\times 10^{6}$ $2.39\times 10^{6}$ 1.24
256 $6.92\times 10^{5}$ $2.44\times 10^{6}$ $2.45\times 10^{6}$ 0.44
Table 2: Model capacity estimates across sequence length $S$, along with error (%).
$V$ Params. Memorized Expected Error
128 $4.21\times 10^{5}$ $1.49\times 10^{6}$ $1.49\times 10^{6}$ 0.36
512 $4.71\times 10^{5}$ $1.71\times 10^{6}$ $1.67\times 10^{6}$ 2.78
1024 $5.36\times 10^{5}$ $1.95\times 10^{6}$ $1.90\times 10^{6}$ 2.70
2048 $6.67\times 10^{5}$ $2.39\times 10^{6}$ $2.36\times 10^{6}$ 1.11
4096 $9.29\times 10^{5}$ $3.13\times 10^{6}$ $3.15\times 10^{6}$ 0.47
Table 3: Model capacity estimates across vocab size $V$, along with error (%).
Figure 12: Model memorization across sequence lengths for a fixed-length dataset. Our predictions of total memorization are accurate, with an average error rate of 1.7%.
Figure 13: Model memorization across vocabulary size for a fixed-length dataset. Our predictions of total memorization are accurate, with an average error rate of 1.8%. Note that, we do not observe a capacity plateau, since increasing $V$ also increases parameters.
Instead of scaling the number of examples in a dataset, we scale model sequence length to adjust the size of a dataset. We use the following measurement for expected memorization of a model:
$\text{mem}(X,L(X))\approx\min(capacity(L),H(X))$
we substitute our previous estimate of $\alpha=3.642$ and ensure to adjust the parameter count for increases due to resizing the model’s embedding matrices. We fix the number of training samples to $4096$ and train a model with $2$ layers and a hidden size of $128$. Results are illustrated in Figure 13 and Table 3. Our predictions of total memorization are accurate, with an average error rate of 1.7% while scaling $S$ and 1.8% when scaling $V$.
A.6 Additional memorization results
Our findings indicate that memorization of text data neatly plateaus near the model capacity just as in the synthetic data case. When the dataset size increases by a factor of $N$, the model divides its memorization between datapoints by an equal amount; the sum of memorization is measured to be constant, presumably at the upper bound of the model’s capacity.
When the dataset is small enough for each model to fit – that is, below the capacity of the smallest model – we observe very similar performance between the models. For larger data sizes we notice an interesting trend: unintended memorization increases with dataset size for to a point, presumably as a model fills its capacity with the available information, and then decreases, as the model replaces sample-level information with more useful, generalizable knowledge. A given model generalizes the most (and memorizes the least information about any individual sample) when the dataset is maximally large.
Figure 14: Train and test losses for different-sized language models trained on synthetic data.
Figure 15: Membership inference attack performance decreases with dataset scale. In the case of uniform synthetic data, membership inference performance never falls below $0.54$.
A.7 Comparison of distributions memorized
Figure 16: Distribution of compression rates for equal-sized transformers ($n_{\text{layer}}=4$, $d_{\text{model}}=128$) trained on $2^{14}$ sequences of equal-length random bitstrings (left) and text (right).
Distribution-level analysis.
Text sequences have very different properties than uniform synthetic bitstrings. We explore how two models of equal capacity spread their memorization across datapoints. We plot a histogram (Figure 16) of train and test compression rates of training data from both synthetic random bitstrings and text. Random training data follows a very normal distribution with a small amount of overlap between train and test compression rates. Text loss is lower on average but more spread out, with low loss on some training points and a long tail of higher losses. There is much more overlap between the train and test loss distributions, which explains why membership inference is more difficult for text data.
Which datapoints are most memorized?
Our distribution-level analysis indicates that unlike in the random-bitstring case, models trained on a large amount of text are able to memorize a small number of datapoints. Prior work has indicated that a large amount of this memorization can be due to duplicated training points (Lee et al., 2022) but our dataset is fully deduplicated so this cannot be an explanation in our case.
Figure 17: Unintended memorization vs. TF-IDF for all training points of a $20M$ param model trained past its capacity on $2^{16}$ sequences of English text. The training documents with rarest words are typically the most memorized.
To quantitatively evaluate the number of rare words per document, we measure the TF-IDF of each training document, plotted vs. unintended memorization in Figure 17. We use the following equation for TF-IDF:
[…]
where $tf(d,\mathcal{D})$ indicates the total number of times word $w$ appears in dataset $\mathcal{D}$. Intuitively, a higher TF-IDF score for document $d$ indicates that $d$ contains more words that are rare in $\mathcal{D}$.
We clearly observe for samples with positive unintended memorization there is a strong correlation between trainset TF-IDF and memorization: examples with more rare words are more memorized. In particular, the sample with highest TF-IDF out of the whole training dataset (a sequence of Japanese words) has the third-highest measured memorization; even though this is just one out of $260,000$ training samples, the model can regurgitate the entire sequence given just a single token ({CJK*}UTF8bkai囚). Out of the top twenty memorized sequences, all but three contain sequences of tokens from other languages (Japanese, Chinese, and Hebrew).
Manual analysis (Table 4) indicates that the most memorized datapoints have extremely rare tokens, typically ones not found in English.
Table 4: Highest TF-IDF training examples from a $20M$ param model trained past its capacity on $2^{16}$ sequences of English text. All of the highest TF-IDF examples are considered memorized, and contain text from non-English languages (Japanese, Chinese, Hebrew, and Greek).
A.8 Scaling law fit
Here we demonstrate the fit of our sigmoidal scaling law to experimental data. We show points in tokens-per-parameter vs. fit in Figure 18. Although the sigmoidal function is slightly simplistic (the points do not perfectly fit the curve) our fit produces estimates within $1-2\%$ of observations.
Figure 18: Our sigmoidal scaling law for membership inference fit to experimental data.
A.9 Proofs
In the section we provide the proofs missing from the main body.
A.10 Proof of Proposition 1
Here we prove Proposition 1
Proof.
we have
$\text{mem}_{U}(X,\hat{\Theta},\Theta)$ $=I(X\mid\Theta,\hat{\Theta})$
$=I((X_{1}\mid\Theta,\dots,X_{n}\mid\Theta),\hat{\Theta}).$
And since the data is sampled i.i.d., all random variables in $\{R_{i}=[X_{i}\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \Theta]\}_{i\in[n]}$ are independent.  [^4] Note that $X_{i}$ themselves are not independent because they are sampled by first sampling an underlying model $\Theta$. However, they are conditionally independent once the underlying model $\Theta$ is given. So we have,
$I((X_{1}\mid\Theta,\dots,X_{n}\mid\Theta),\hat{\Theta})\geq\sum_{i\in[n]}I(X_{i}\mid\Theta,\hat{\Theta})$
which implies
$\text{mem}_{U}(X,\hat{\Theta},\Theta)\geq\sum_{i\in[n]}{\text{mem}_{U}(X_{i},\hat{\Theta}},\Theta).$
On the other hand, we have
$\text{mem}_{U}(X,\hat{\Theta},\Theta)$ $=I(X\mid\Theta,\hat{\Theta})$
$=H(\hat{\Theta}-H(\hat{\Theta}\mid(X\mid\Theta))$
$\leq H(\hat{\Theta})$
A.11 Proof of Proposition 4
Proof.
We first state a Lemma about connection between algorithmic (kolmogorov) mutual information and mutual information.
Lemma 6.
[Theorem 3.6 in Grunwald & Vitányi (2004)] Assume $(X,Y)$ be a pair of joint random variables. Let $f$ be the density function, $f(x,y)=\Pr[(X,Y)=(x,y)].$ Then we have
$I(X,Y)-H_{K}(f)$ $\leq\operatorname*{E}_{(x,y)\sim(X,Y)}[I_{K}(x,y)]$
$\leq I(X,Y)+2H_{K}(f).$
Now we use this lemma to prove the statement of the Proposition. Let $f$ be a the density function for the joint distribution $(X_{i}\mid\theta,\hat{\Theta})$. That is $f_{i}(x_{i},\hat{\theta})=\Pr[X_{i}=x_{i}\mid\theta\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\hat{\Theta}=\hat{\theta}]$. Note that this function is independent of $n$ and $\theta$. By definition we have
$\text{mem}_{U}(X_{i},\hat{\Theta},\theta)=I(X_{i}\mid\theta,\hat{\Theta}).$
Now using Lemma 6 we have
$I(X_{i}\mid\theta,\hat{\Theta})-H_{K}(f)$ $\leq\operatorname*{E}_{x_{i}\sim X_{i}\mid\theta}[I_{K}(x_{i},\hat{\theta})]$
$\leq I(X_{i}\mid\theta,\hat{\Theta})+2H_{K}(f).$
and this concludes the statement of Proposition by setting $\epsilon=2H_{K}(f)$
A.12 Limitations
Our efforts to measure language model memorization come from a line of recent research to discover whether models have analyzed certain texts, and if so, how much. However, our main experimental contributions relate to the practice of training and evaluating language models, including a new perspective on the phenomenon of grokking (Nakkiran et al., 2019) and a new measurement of capacity. Our results are specific to the environment proposed and do not necessarily generalize to other datasets, architectures, or training setups.